1. A wagon weighs 12.0 lb, and the coefficient of rolling friction between the wheels and the floor is 0.140. What horizontal pull is need to keep the wagon rolling in a straight line at 2.00 ft/s?

2. A wagon weighs 50.0 N, and the coefficient of rolling friction between the wheels and the level floor is 0.150. What horizontal pull is need to keep the wagon rolling uniformly?
Answer #1

1)

We need to use Newton’s 2nd law of motion here,

Fnet = m*a

Fpull – Ffriction = m*a

Fpull – C*Fn = m*a

Where C= rolling coefficient , and Fn= normal force

Fpull – C*mg = m*a -------------------(1)

a=0 Ft/s^2 since velocity v=2.0ft/s is constant.

Plugging given data,

Fpull – 0.140*12.0*32.17 = 1.20*0

Fpull = 0.140*12.0*32.17

Fpull = 54.05 lb

b)

mg=50.0 N, C=0.150 and a= 0 m/s^2

Plugging given data in eqn (1),

Fpull – C*mg = m*a

Fpull – 0.150*50=(50/9.8)*0

Fpull = 0.150*50

Fpull = 7.5 N